OpenStudy (anonymous):

Which of the following represents 3x - 5y + 10 = 0 written in slope-intercept form?

OpenStudy (texaschic101):

slope intercept form : y = mx + b 3x - 5y + 10 = 0 -5y = -3x - 10 y = 3/5x + 2